Aj749418:

Left on together, the cold and hot water faucets of a certain bathtub take 4 minutes to fill the tub. If it takes the hot water faucet 14 minutes to fill the tub by itself, how long will it take the cold water faucet to fill the tub on its own?

AZ:

That's a similar question but not the same numbers so don't get the numbers confused. If you turn on the cold and hot water faucets, it takes 4 minutes to fill the tub. If you only turn on the hot water faucet, it takes 14 minutes to fill the tub. If you only turn on the cold water faucet, it takes `x` minutes to fill the tub. For the equation we're about to come up with, we write it as 1/x because that's how much of the tub will get filled in 1 minute. And so for the entire tub to get filled, it would take 'x' minutes. The question asks us to find out how long it takes to fill the tub with cold water so basically finding 'x'. Our equation will be \(\dfrac{1}{4} = \dfrac{1}{14} + \dfrac{1}{x}\) Now all we have to do is solve for x. Can you do that?

Aj749418:

Yep the answer is 5.6. Thanks!

AZ:

Of course! Glad I could help :)